What number subtracted from each 71 and 58 will result in two perfect square?

Answer by trsomas23@gmail.com(17)

The greatest number less than 58, which is a perfect square is 49 because 49 is square of 7.
71 - 49 = 22
71 - 22 = 49
58 - 22 = 36, which is also a perfect square.
Answer: 22